• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Kumbartzki

    Log entry time 08:04:34 on July28,2006

    Entry number 175665

    keyword=Owl shift summary

    Hall was in controlled access until 1:08
    Setup for H elastic kinematics Point 1 0.5577 GeV/c
    Started with optics.
    good runs are 1757 - 1759 (before the raster was not working, also the prescaler ps3 was too high). At end lost beam.
    Switched to LH2 and started run 1760 at 2:37. Take data with 8uA since trigger rate is only 1k. Deadtime is about 16%.
    After 2 hours the file split, but we stopped and started a new run.
    The split file is very short.
    6:45 changed to kin point 2 (0.5467 GeV/c). The trigger rate was now more than 3 k, reduced beam to 4 uA which gives 24% dead time and 1.9 k rate
    started run 1763